a stone is dropped from the top of a tower and travels 24.5m in the last second of its journey. the height of the tower is


(a)44.1


(b)49


(c)78.4


(d)72

Grade:

1 Answers

sagar garg
11 years ago

Displacement in n th second=u+1/2*a*(2n-1)

                                  -24.5=0+1/2*(-9.8)*(2n-1)

                                   2n-1=5

                                    n=3

Time taken by body to cover height of tower=3seconds

Sice u=0;Displacement =1/2*g*t2

                            Height=1/2*9.8*(3)2  

                                    =44.1

answer:(a)